In a certain Algebra 2 class of 30 students, 7 of them play basketball and 15 of them play baseball. There are 10 students who play neither sport. What is the probability that a student chosen randomly from the class plays basketball or baseball?



1
Expert's answer
2022-02-28T12:14:56-0500
"P(baseball\ or\ basketball)=\dfrac{30-10}{30}=\dfrac{2}{3}"

The probability that a student chosen randomly from the class plays basketball or baseball is 2/3.